A lot consists of 10 articles, 4 with minor defects, and 3 with major defects. Two articles are [05] chosen at random from the lot without replacement. Compute the probability by using basic probability concepts, both articles are good.
1
Expert's answer
2020-12-18T14:50:51-0500

Number of good articles = 10-4-3 = 3

P(both articles are good) = P(first article is good)*P(second article is good if first article is good) =

3/10*2/9 = 1/15 = 0.0667